INTRAMOLECULAR PROCESSES OF EXITED STATE OF SALICYLIDENEANILINE MOLECULAR CLASS

\underline{D. G. TOVBIN}, F. E. GOSTEV, A. N. PETRUKHIN, A. A. TITOV, V. A. NADTOCHENKO, and O. M. SARKISOV, Laser Spectroscopy Laboratory, N.Semenov Institute of Chemical Physics RAS, Moscow, Russia 119991; A.V. METELITSA, V.I. MINKIN, Institute of Physical and Organic Chemistry, Rostov State University, Rostov on Don, Russia, 344090.

The intramolecular light-induced H+ transfer was studied in the family of salicylideneaniline like (SA) moleculs by femtosecond laser absorption spectroscopy with time resolution 80 fs^{a}. The characteristic times of the proton transfer step and futher relaxation process were determined. The evidences of two twisted intramolecular charge transfer structures were obtained. The key role of acoplanarization at the photochromic properties is discussed.

aGostev F.E., at al. Russian Chemical Bulletin, 2004 in press.